### v1

@@ 6.2. Linear multislicing
$\max_{i=1,\dots q}{\mathcal{N}}_{\rho}(\pi_{F_{i}}A^{\prime\prime})<\rho^{-\alpha k-\varepsilon}.$
@@
$\begin{split}{\mathcal{E}}:=\{\,F\in\operatorname{Gr}(\mathbb{R}^{D},k)\,:\,\exists A^{\prime}\subseteq A\,\,&\text{ with }\,\,{\mathcal{N}}_{\rho}(A^{\prime})\geq\rho^{\varepsilon}{\mathcal{N}}_{\rho}(A)\\ &\text{ and }\,\,{\mathcal{N}}_{\rho}(\pi_{F}A^{\prime})<\rho^{C\varepsilon}{\mathcal{N}}_{\rho}(A)^{\frac{k}{D}}\}\end{split}$
@@
The proof is similar to that of Lemma˜6.12, using the subcritical projection theorem [6, Proposition A.2] instead of the supercritical projection theorem. ∎
We now combine Lemmas 6.12, 6.13 into a multislicing estimate. We place ourselves in $\mathbb{R}^{D}$ where $D\geq 3$. We consider $d_{1},d_{2}\in\mathbb{N}$ such that $1\leq d_{1}<d_{2}<D$ and ${\bf t}=(t_{1},t_{2},t_{3})\in\mathbb{R}^{3}$ such that $0\leq t_{1}<t_{2}<t_{3}\leq 1$. We set $\mathcal{F}$} the collection of couples $\mathscr{V}=(V_{1},V_{2})$ where $V_{i}\in\operatorname{Gr}(\mathbb{R}^{d},d_{i})$ for $i=1,2$ and $V_{1}\subseteq V_{2}$. Given $\mathscr{V}\in{\mathcal{F}}$ and $\rho\in(0,1)$, we set
@@
Let $\nu$ be a Borel measure on $B^{\mathbb{R}^{D}}_{1}$ such that for some $\alpha\in[c,1-c]$, for all $v\in\mathbb{R}^{D}$, and $r\geq\rho$, we have
@@
Proof.
Lemmas 6.12, 6.13 guarantee that the random projectors $\pi_{||V_{1}}$ and $\pi_{||V_{2}}$ where $\mathscr{V}\sim\Xi$ satisfy respectively subcritical and supercritical estimates. Those can be combined as in the original paper [6, Section 2] into the above result. More formally, the deduction is a direct consequence of [5, Theorem 3.4].∎
@@ 6.3. Linearizing charts
Recall $X=G/\Lambda$ where $G=\operatorname{SL}_{d+1}(\mathbb{R})$ and $\Lambda$ is a fixed arbitrary lattice. We define on $X$ a covering of linearizing charts which do not deform balls much, and most importantly send any $g$-translate $gB_{r}x$ ($g\in G,r>0,x\in X$) to an additive translate of the box $\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}$ provided $\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}$ is not too distorted and lives at a suitable scale. We point out that contrary to our previous work [7] where $d=1$, those charts live at a microscopic scale. This linearizing scheme is extracted from [5, Lemma 6.3], which is itself inspired by Shmerkin [51].
Lemma 6.15 ([5]).
Let $0<\delta\leq\eta\lll 1$. There exists a measurable map $\varphi:\{\operatorname{inj}\geq\eta\}\rightarrow B^{\mathfrak{g}}_{1}$ satisfying the following.
@@
    For every $r\in(0,\eta)$, $v\in\mathfrak{g}$, the preimage $\varphi^{-1}(B^{\mathfrak{g}}_{r}+v)$ is covered by $O(1)$ many balls $(B_{r}x)_{x\in X}$
@@
    For every $r\in(0,\eta)$, $g\in G$ such that $B^{\mathfrak{g}}_{\delta^{2}}\subseteq\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}\subseteq B^{\mathfrak{g}}_{\delta}$, and $x\in X$, the translate $gB_{r}x\cap\{\operatorname{inj}\geq\eta\}$ is covered by $O(1)$ many preimages of boxes $(\varphi^{-1}(\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}+v))_{v\in\mathfrak{g}}$.
In this lemma, $\eta$ controls the region of $X$ which is linearized and the maximum scale at which linearization occurs. On the other hand, $\delta$ controls the distortion allowed on $G$-translates of balls $gB_{r}x$ to be well represented by additive translates of boxes $\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}+v$ via the linearization.

### v2

@@ 7.2. Linear multislicing
$\max_{i=1\dots q}{\mathcal{N}}_{\rho}\bigl(\pi_{F_{i}}S\bigr)\geq\rho^{O_{D,c}(\varepsilon_{1})}\lvert S\rvert^{1-1/q}{\mathcal{N}}_{\rho}(\pi_{\bigcap_{i}F_{i}}S^{\prime})^{1/q}.$ (7.20)
Taking $S$ to be a not too small subset of $A$, we will use (7.20) to obtain an explicit lower bound on $\max_{i=1,\dots,q}{\mathcal{N}}_{\rho}(\pi_{F_{i}}S)$, see (7.22). As a lower bound on $\lvert S\rvert^{q-1}$ comes directly from the assumption (7.18), we focus on ${\mathcal{N}}_{\rho}(\pi_{\bigcap_{i}F_{i}}S^{\prime})$.
Let $\Upsilon$ denote the restriction of $\Xi^{\otimes q}$ to $\mathcal{G}^{q}\smallsetminus{\mathcal{E}}_{1}$, renormalised into a probability measure. Note that, by definition and the observation that
$\bigl\lVert(F_{1}^{\perp}\oplus\dotsb\oplus F_{q}^{\perp})\wedge W\bigr\rVert\geq\bigl\lVert F_{1}^{\perp}\wedge\dotsm\wedge F_{q}^{\perp}\wedge W\bigr\rVert,$
the random Dq(Dk)D-q(D-k)-plane $(\bigcap_{i}F_{i})_{{\underline{F}}\sim\Upsilon}$ satisfies the non-concentration condition $\operatorname{(MNC)}^{\perp}$ with parameters $(\rho,\rho^{-\varepsilon-\varepsilon_{1}},c)$ and $q=1$, allowing us to apply the $q=1$ case. Therefore, provided that $\varepsilon+\varepsilon_{1},\varepsilon_{2}\lll_{D,c}1$ and $\rho\lll_{D,c}1$, there is an event ${\mathcal{E}}_{2}\subseteq\mathcal{G}^{q}$ such that $\Upsilon({\mathcal{E}}_{2})\leq\rho^{\varepsilon_{2}}$, and for all ${\underline{F}}\in\mathcal{G}^{q}\smallsetminus{\mathcal{E}}_{2}$, for all $A^{\prime}\subseteq A$ with $\lvert A^{\prime}\rvert\geq\rho^{\varepsilon_{2}}\lvert A\rvert$, we have
${\mathcal{N}}_{\rho}\bigl(\pi_{\bigcap_{i}F_{i}}A^{\prime}\bigr)\geq\rho^{-\alpha(D-q(D-k))-\varepsilon_{2}}.$ (7.21)
Combining (7.20), (7.18) and (7.21), we obtain that for all ${\underline{F}}\in\mathcal{G}^{q}\smallsetminus({\mathcal{E}}_{1}\cup{\mathcal{E}}_{2})$, and $A^{\prime\prime}\subseteq A$ with $\lvert A^{\prime\prime}\rvert\geq\rho^{\varepsilon_{1}}\lvert A\rvert$, we have
$\max_{i=1,\dots,q}{\mathcal{N}}_{\rho}\bigl(\pi_{F_{i}}A^{\prime\prime}\bigr)\geq\rho^{-\alpha k-\frac{1}{2q}\varepsilon_{2}},$ (7.22)
provided that $\varepsilon\leq\varepsilon_{1}\lll_{D,c}\varepsilon_{2}\lll_{D,c}1$.
To conclude the proof, we argue by contradiction, assuming $\Xi({\mathcal{E}})>\rho^{\varepsilon}$. For each $F\in{\mathcal{E}}$, let $A^{\prime}_{F}\subseteq A$ be such that $\lvert A^{\prime}_{F}\rvert\geq\rho^{\varepsilon}\lvert A\rvert$ and ${\mathcal{N}}_{\rho}(\pi_{F}A^{\prime}_{F})<\rho^{-\alpha k-\varepsilon}$. By [29, Lemma 19] – a Fubini argument – applied to $q$ independent copies of $A^{\prime}_{F}$, where $F$ is distributed according to $\Xi_{|{\mathcal{E}}}$, these independent copies are likely to intersect in a rather large subset:
${\mathcal{I}}:=\bigl\{\,{\underline{F}}\in{\mathcal{E}}^{q}\,:\,\lvert A^{\prime}_{F_{1}}\cap\dots\cap A^{\prime}_{F_{q}}\rvert\leq\rho^{2q\varepsilon}\lvert A\rvert\,\bigr\}\qquad\text{satisfies}\qquad\Xi^{\otimes q}({\mathcal{I}})\geq\rho^{4q\varepsilon}\Xi({\mathcal{E}})^{q}\geq\rho^{5q\varepsilon}.$
It follows that $\Xi^{\otimes q}\bigl({\mathcal{I}}\smallsetminus({\mathcal{E}}_{1}\cup{\mathcal{E}}_{2})\bigr)\geq\rho^{5q\varepsilon}-\rho^{\varepsilon_{1}}-\rho^{\varepsilon_{2}}$ which is strictly positive provided that $\varepsilon\lll_{D}\varepsilon_{1}\leq\varepsilon_{2}$ and $\rho\lll_{\varepsilon}1$. In particular, we may consider ${\underline{F}}\in{\mathcal{I}}\smallsetminus({\mathcal{E}}_{1}\cup{\mathcal{E}}_{2})$. Setting $A^{\prime\prime}=A^{\prime}_{F_{1}}\cap\dots\cap A^{\prime}_{F_{q}}$, we have $\lvert A^{\prime\prime}\rvert\geq\rho^{2q\varepsilon}\lvert A\rvert$ while the inclusions $A^{\prime\prime}\subseteq A^{\prime}_{F_{i}}$ yield
$\max_{i=1,\dotsc,q}{\mathcal{N}}_{\rho}\bigl(\pi_{F_{i}}A^{\prime\prime}\bigr)<\rho^{-\alpha k-\varepsilon}.$
@@
$\begin{split}{\mathcal{E}}:=\bigl\{\,F\in\operatorname{Gr}(\mathbb{R}^{D},k)\,:\,\exists A^{\prime}\subseteq A\,\,&\text{ with }\,\,{\mathcal{N}}_{\rho}(A^{\prime})\geq\rho^{\varepsilon}{\mathcal{N}}_{\rho}(A)\\ &\text{ and }\,\,{\mathcal{N}}_{\rho}(\pi_{F}A^{\prime})<\rho^{C\varepsilon}{\mathcal{N}}_{\rho}(A)^{\frac{k}{D}}\,\bigr\}\end{split}$
@@
Remark.
There is no condition on $\varepsilon$ in the statement. However, it is only meaningful for $C\varepsilon\leq D$, which requires in particular $\varepsilon\lll_{D,c}1$. Indeed, if $C\varepsilon>D$ and $\rho\lll_{D}1$, then $\mathcal{E}$$\varnothing$.
@@
The case $q=1$ is the subcritical projection theorem as stated in [4, Proposition A.2]. The general case can be deduced similarly to the proof of Lemma˜7.12, but using this time the subcritical projection theorem instead of the supercritical projection theorem. ∎
We now combine Lemmas 7.12, 7.13 into a multislicing estimate. We place ourselves in $\mathbb{R}^{D}$ where $D\geq 3$. We consider $d_{1},d_{2}\in\mathbb{N}$ such that $1\leq d_{1}<d_{2}<D$ and ${\bf t}=(t_{1},t_{2},t_{3})\in\mathbb{R}^{3}$ such that $0\leq t_{1}<t_{2}<t_{3}\leq 1$. We denote by $\mathcal{F}$} the collection of pairs $\mathscr{V}=(V_{1},V_{2})$ where $V_{i}\in\operatorname{Gr}(\mathbb{R}^{D},d_{i})$ for $i=1,2$ and $V_{1}\subseteq V_{2}$. Given $\mathscr{V}\in{\mathcal{F}}$ and $\rho\in(0,1)$, we set
@@
Let $\nu$ be a Borel measure on $B^{\mathbb{R}^{D}}_{1}$ of mass at most $\rho^{-\varepsilon}$, and such that for some $\alpha\in[c,1-c]$, for all $v\in\mathbb{R}^{D}$, all $r\in[\rho^{t_{3}},\rho^{t_{1}}]$, we have
@@
The proof is done by combining Lemmas 7.12, 7.13 with the line of reasoning of [4, Section 2]. Alternatively, those two lemmas can also be plugged into the more advanced multislicing machinery [5, Theorem 3.4] to obtain Proposition˜7.14 as a formal consequence. For the reader’s convenience, we sketch below how to deduce Proposition˜7.14 from Lemmas 7.12, 7.13.
Sketch of proof.
In this proof, we allow the implied constant in the notations $\ll$ and $O(\cdot)$ to depend on $D$ and $\bf t$}. We may [^6] Indeed, if we establish the proposition for a pair $(\varepsilon,\rho)$ then it is automatically valid for $(\varepsilon^{\prime},\rho)$ with $\varepsilon^{\prime}\in(0,\varepsilon)$, because when passing from $\varepsilon$ to $\varepsilon^{\prime}$, assumptions get stronger and the conclusion gets weaker. also allow $\rho$ to be small enough depending on $\varepsilon$ (not only $D,{\bf t},c$).
Note that if $\nu(\mathbb{R}^{D})\leq\rho^{\varepsilon}$, then we can simply take $A_{\mathscr{V}}=\varnothing$ and the result is trivial. Thus, we may assume $\nu(\mathbb{R}^{D})\in[\rho^{\varepsilon},\rho^{-\varepsilon}]$, and after a renormalisation, reduce to the case where $\nu(\mathbb{R}^{D})=1$.
We further reduce to the situation where $\nu$ is the uniform probability measure on a finite set with a regular-tree structure. For $i=1,2,3$, set ${\mathcal{Q}}_{i}={\mathcal{D}}_{\rho^{t_{i}}}$ the tiling of $\mathbb{R}^{D}$ by cubes of side length $\rho^{t_{i}}$, in particular ${\mathcal{Q}}_{1}\prec{\mathcal{Q}}_{2}\prec{\mathcal{Q}}_{3}$. By invoking [5, Lemma A.2] (which only relies on dyadic pigeonholing) we obtain pairwise disjoint ${\mathcal{Q}}_{3}$-measurable sets $A_{k}\subseteq\mathbb{R}^{D}$ indexed by a set $\mathcal{K}$} of cardinality $\lvert{\mathcal{K}}\rvert\ll\lvert\log\rho\rvert^{O(1)}$ such that $\nu\bigl(\mathbb{R}^{D}\smallsetminus\bigcup_{k\in K}A_{k}\bigr)\leq\rho^{2\varepsilon}$ and satisfying for each $k\in{\mathcal{K}}$:
  • $A_{k}$ is regular with respect to the filtration ${\mathcal{Q}}_{1}\prec{\mathcal{Q}}_{2}\prec{\mathcal{Q}}_{3}$, in the sense of [5, §A.2];
  • $\nu(A_{k})\geq\rho^{8\varepsilon}$;
  • for every ${\mathcal{Q}}_{3}$-measurable subset $S\subseteq A_{k}$, we have
    $\frac{\nu(S)}{\nu(A_{k})}\simeq\frac{{\mathcal{N}}_{\rho^{t_{3}}}(S)}{{\mathcal{N}}_{\rho^{t_{3}}}(A_{k})}.$
The bounds $\nu\bigl(\mathbb{R}^{D}\smallsetminus\bigcup_{k\in K}A_{k}\bigr)\leq\rho^{2\varepsilon}$ and $\lvert{\mathcal{K}}\rvert\ll\lvert\log\rho\rvert^{O(1)}$ justify it is sufficient to prove the conclusion of Proposition˜7.14 for each measure $\nu_{|A_{k}}/\nu(A_{k})$. By discretizing at scale $\rho^{t_{3}}$, this in turn reduces to the case of the measure $\frac{1}{{\mathcal{N}}_{\rho^{t_{3}}}(A_{k})}\sum_{Q\in{\mathcal{Q}}_{3}}\delta_{x_{Q}}$, where $x_{Q}$ denotes the center of the $\delta^{t_{3}}$-cell $Q$. Note these measures satisfy the non-concentration conditions required in Proposition˜7.14 (with $9\varepsilon$ instead of $\varepsilon$). We are thus reduced to the case where $\nu$ is the uniform probability measure on a finite set $A=\operatorname{supp}\nu$, which is regular for ${\mathcal{Q}}_{1}\prec{\mathcal{Q}}_{2}\prec{\mathcal{Q}}_{3}$, and intersects each cell of ${\mathcal{Q}}_{3}$ in at most one point.
Let ${\mathcal{P}}^{\mathscr{V}}$ denote a tiling of $\mathbb{R}^{D}$ by translates of a prism of shape comparable to $B^{\mathscr{V}}_{\rho^{\bf t}}$. Let ${\mathcal{P}}^{\mathscr{V}}_{\mathrm{bad}}$ denote the subset of tiles $P\in{\mathcal{P}}^{\mathscr{V}}$ such that $\nu(P)>2^{-D}\operatorname{Leb}\bigl(B^{\mathscr{V}}_{\rho^{\bf t}}\bigr)^{\alpha+\varepsilon}$. It suffices to show that $\Xi({\mathcal{E}})\leq\rho^{\varepsilon}$ for
${\mathcal{E}}:=\bigl\{\,\mathscr{V}\in{\mathcal{F}}\,:\,\nu(\cup{\mathcal{P}}^{\mathscr{V}}_{\mathrm{bad}})>\rho^{\varepsilon}\,\bigr\}.$
Let $\mathscr{V}\in{\mathcal{E}}$ be arbitrary. Let $E$ denote the smallest ${\mathcal{Q}}_{3}$-measurable set containing $\cup{\mathcal{P}}^{\mathscr{V}}_{\mathrm{bad}}$. On the one hand, using the structure of $\nu$ and $\nu(E)>\rho^{\varepsilon}$, we find
${\mathcal{N}}_{\rho^{t_{3}}}(E)\gg\rho^{\varepsilon}{\mathcal{N}}_{\rho^{t_{3}}}(A).$ (7.23)
On the other hand, $E$ is contained in the $\rho^{t_{3}}$-neighborhood of $\cup{\mathcal{P}}^{\mathscr{V}}_{\mathrm{bad}}$ and $\rho^{t_{3}}$ is the length of the shortest side of tiles in ${\mathcal{P}}^{\mathscr{V}}$. Hence,
${\mathcal{N}}_{{\mathcal{P}}^{\mathscr{V}}}(E)\ll{\mathcal{N}}_{{\mathcal{P}}^{\mathscr{V}}}(\cup{\mathcal{P}}^{\mathscr{V}}_{\mathrm{bad}})\leq\operatorname{Leb}(P)^{-\alpha-2\varepsilon}$ (7.24)
where the last bound uses the definition of ${\mathcal{P}}^{\mathscr{V}}_{\mathrm{bad}}$ and that $\nu(\mathbb{R}^{D})\leq 1$.
Next, consider two tiling ${\mathcal{S}}^{\mathscr{V}}$ and ${\mathcal{R}}^{\mathscr{V}}$ of $\mathbb{R}^{D}$ consisting respectively of prisms of shape approximately $B^{\mathscr{V}}_{\rho^{(t_{1},t_{2},t_{2})}}$ and $B^{\mathscr{V}}_{\rho^{(t_{2},t_{2},t_{3})}}$, in particular ${\mathcal{S}}^{\mathscr{V}}\overset{O(1)}{\prec}{\mathcal{P}}^{\mathscr{V}},{\mathcal{Q}}_{2}$, and ${\mathcal{R}}^{\mathscr{V}}\overset{O(1)}{\simeq}{\mathcal{P}}^{\mathscr{V}}\vee{\mathcal{Q}}_{2}$. To anticipate a technical detail, we replace $E$ by a large subset which is moreover regular with respect to ${\mathcal{Q}}_{2}\overset{O(1)}{\prec}{\mathcal{R}}^{\mathscr{V}}\overset{O(1)}{\prec}{\mathcal{Q}}_{3}$. This can be done (using e.g. [4, Lemma 2.5]) while keeping trivially (7.24) and slightly weakening (7.23) to
${\mathcal{N}}_{\rho^{t_{3}}}(E)\geq\rho^{2\varepsilon}{\mathcal{N}}_{\rho^{t_{3}}}(A).$ (7.25)
By our choices for ${\mathcal{S}}^{\mathscr{V}},{\mathcal{R}}^{\mathscr{V}}$, Lemma˜7.10 yields the existence of a subset $E^{\prime}\subseteq E$ such that
${\mathcal{N}}_{{\mathcal{R}}^{\mathscr{V}}}(E^{\prime})\gg{\mathcal{N}}_{{\mathcal{R}}^{\mathscr{V}}}(E)$ (7.26)
and
${\mathcal{N}}_{{\mathcal{R}}^{\mathscr{V}}}(E){\mathcal{N}}_{{\mathcal{S}}^{\mathscr{V}}}(E^{\prime})\ll{\mathcal{N}}_{{\mathcal{P}}^{\mathscr{V}}}(E){\mathcal{N}}_{\rho^{t_{2}}}(E),$ (7.27)
Using $E\subseteq A$ and (7.24) , this yields the (crucial) inequality
${\mathcal{N}}_{{\mathcal{R}}^{\mathscr{V}}}(E){\mathcal{N}}_{{\mathcal{S}}^{\mathscr{V}}}(E^{\prime})\ll\operatorname{Leb}(P)^{-\alpha-2\varepsilon}{\mathcal{N}}_{\rho^{t_{2}}}(A).$ (7.28)
We will now apply the projection estimates established in Lemma˜7.12, Lemma˜7.13 to bound from below ${\mathcal{N}}_{{\mathcal{R}}^{\mathscr{V}}}(E)$ and ${\mathcal{N}}_{{\mathcal{S}}^{\mathscr{V}}}(E^{\prime})$ for most $\mathscr{V}$ selected by $\Xi$, and ultimately get a contradiction with (7.28).
Let us bound from below ${\mathcal{N}}_{{\mathcal{S}}^{\mathscr{V}}}(E^{\prime})$. It essentially represents the covering number of $E^{\prime}$ by translates of the box $B^{\mathscr{V}}_{\rho^{(t_{1},t_{2},t_{2})}}$. This box can be seen, locally at scale $\rho^{t_{1}}$, as the preimages of a ball of radius $\rho^{t_{2}}$ by the orthogonal projector of kernel $V_{1}$. This motivates the application of a projection theorem. To do so, we zoom into each $Q\in{\mathcal{Q}}_{1}(A)$ by introducing $A^{Q}$ the $\times\rho^{-t_{1}}$-dilation of $A\cap Q$, and we apply one of the projection estimates to the set $A^{Q}$ at scale $\rho^{-t_{1}+t_{2}}$ and the random projector $(\pi_{||V_{1}})_{\mathscr{V}\sim\Xi}$. In principle, this yields a lower bound for ${\mathcal{N}}_{\rho^{-t_{1}+t_{2}}}(\rho^{-t_{1}}(E^{\prime}\cap Q))\simeq{\mathcal{N}}_{{\mathcal{S}}^{\mathscr{V}}}(E^{\prime}\cap Q)$ because $E^{\prime}\cap Q$ is a large subset of $A\cap Q$ for most $Q$ (due to ${\mathcal{N}}_{\rho^{t_{2}}}(E^{\prime})\gg\rho^{3\varepsilon}{\mathcal{N}}_{\rho^{t_{2}}}(A)$ which stems from (7.25), (7.26) and the regularity properties of $A$, $E$). Summing over $Q$, we obtain in the end a lower bound on ${\mathcal{N}}_{{\mathcal{S}}^{\mathscr{V}}}(E^{\prime})$. Indeed, if ${\mathcal{N}}_{\rho^{t_{i}}}(A)>\rho^{-t_{i}D\alpha-10D\varepsilon}$ holds for some $i\in\{1,2\}$, we apply the subcritical version Lemma˜7.13 to each $A^{Q}$, and ultimately get
${\mathcal{N}}_{{\mathcal{S}}^{\mathscr{V}}}(E^{\prime})$ $\geq\rho^{4\varepsilon}{\mathcal{N}}_{\rho^{t_{1}}}(A)^{d_{1}/D}{\mathcal{N}}_{\rho^{t_{2}}}(A)^{D-d_{1}/D}$
$\geq\rho^{-\alpha(t_{1}d_{1}+t_{2}(d_{2}-d_{1}))-5\varepsilon}{\mathcal{N}}_{\rho^{t_{2}}}(A)^{D-d_{2}/D}$ (7.29)
where the second inequality also uses the Frostman-type condition on $\nu$. If on the other hand ${\mathcal{N}}_{\rho^{t_{i}}}(A)\leq\rho^{-t_{i}D\alpha-10D\varepsilon}$ for both $i=1,2$, we are in a position to apply the supercritical projection theorem Lemma˜7.12 to still obtain (7.29). Indeed, in this case, the Frostman-type condition on $\nu$ together with the regularity properties of $A$ imply that for every ball $B\subseteq Q$ of radius $r\in[\rho^{t_{2}},\rho^{t_{1}}]$,
${\mathcal{N}}_{\rho^{t_{2}}}(B\cap A)\leq\rho^{-20D\varepsilon}\left(\frac{r}{\rho^{t_{1}}}\right)^{D\alpha}{\mathcal{N}}_{\rho^{t_{2}}}(Q\cap A).$
This is precisely the required non-concentration needed to apply Lemma˜7.12 to $A^{Q}$. In the end, we have justified (7.29) regardless of the value of $({\mathcal{N}}_{\rho^{t_{i}}}(A))_{i=1,2}$.
To bound from below ${\mathcal{N}}_{{\mathcal{R}}^{\mathscr{V}}}(E)$, we proceed similarly, by zooming into each $Q\in{\mathcal{Q}}_{2}(A)$ and applying Lemma˜7.13 to the $\times\rho^{-t_{2}}$-dilation of $A\cap Q$ and the random projector $(\pi_{||V_{2}})_{\mathscr{V}\sim\Xi}$. We obtain the subcritical estimate
${\mathcal{N}}_{{\mathcal{R}}^{\mathscr{V}}}(E)\geq\rho^{-\alpha t_{3}(D-d_{2})+2\varepsilon}{\mathcal{N}}_{\rho^{t_{2}}}(A)^{d_{2}/D}.$ (7.30)
Now (7.28), (7.29), (7.30) are in contradiction, which concludes the proof. ∎
@@ 7.3. Linearizing charts
Recall $X=G/\Lambda$ where $G=\operatorname{SL}_{d+1}(\mathbb{R})$ and $\Lambda$ is a fixed arbitrary lattice. We define on $X$ a covering of linearizing charts which do not deform balls much, and most importantly send any $g$-translate $gB_{r}x$ ($g\in G,r>0,x\in X$) to an additive translate of the box $\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}$ provided that $\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}$ is not too distorted and lives at a suitable scale. We point out that these charts live at a microscopic scale, contrary to those used in our previous work [3] about the case $d=1$. This linearizing scheme is extracted from [5, Lemma 6.3], which is itself inspired by Shmerkin [59].
Lemma 7.15 ([5, Lemma 6.3]).
Let $0<\delta\lll 1$. There exists a measurable map $\varphi:\{\operatorname{inj}\geq\delta\}\rightarrow B^{\mathfrak{g}}_{1}$ satisfying the following.
@@
    For every $r\in(0,\delta)$, $v\in\mathfrak{g}$, the preimage $\varphi^{-1}(B^{\mathfrak{g}}_{r}+v)$ is covered by $O(1)$ many balls $(B_{r}x)_{x\in X}$
@@
    For every $r\in(0,\delta)$, $g\in G$ such that $B^{\mathfrak{g}}_{\delta^{2}}\subseteq\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}\subseteq B^{\mathfrak{g}}_{\delta}$, and $x\in X$, the translate $gB_{r}x\cap\{\operatorname{inj}\geq\delta\}$ is covered by $O(1)$ many preimages of boxes $(\varphi^{-1}(\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}+v))_{v\in\mathfrak{g}}$.
In this lemma, the inequality $B^{\mathfrak{g}}_{\delta^{2}}\subseteq\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}\subseteq B^{\mathfrak{g}}_{\delta}$ controls the distortion allowed on $G$-translates of balls $gB_{r}x$ to be well represented by additive translates of boxes $\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}+v$ via the linearization.